OpenStudy (anonymous):

I know that this i a parabola, but I need help putting it into vertex form: y=2x^2-3x+4

OpenStudy (anonymous):

\[y=2x ^{2}-3x+4\]

OpenStudy (ash2326):

sorry vertex is (3/4, 23/4)

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

y= 2( x^2-3/2 x+2) y=2(x^2-3/2 x+9/16-9/16+2) y=2(x^2-1.5x+9/16)-9/8+4 y=2(x-3/4)^2+23/8

OpenStudy (ash2326):

sorry i made a mistake it'll be (3/4, 23/8)

OpenStudy (anonymous):

thanks you guys :)